# UniChange: Unifying Change Detection with Multimodal Large Language Model

### v1

@@ 3.3 Training Strategy
Our method is trained end-to-end to jointly optimize UniChange. The overall objective function, $\mathcal{L}_{total}$, is defined as a weighted summation of all contributing loss terms:
$\mathcal{L}_{total}=\lambda_{txt}\mathcal{L}_{txt}+\mathcal{L}_{mask},$ (11)
the comprehensive mask loss, $\mathcal{L}_{mask}$, is designed to enforce pixel-level accuracy across the unified change detection task and is defined as the summation of four distinct components:
$\mathcal{L}_{mask}=\lambda_{BCE}\mathcal{L}_{BCE}+\lambda_{Dice}\mathcal{L}_{Dice}+\lambda_{SS}\mathcal{L}_{SS}+\lambda_{CS}\mathcal{L}_{CS},$
in Eq. 12, $\lambda_{txt}$, $\lambda_{BCE}$, $\lambda_{Dice}$, $\lambda_{SS}$, and $\lambda_{CS}$ serve as the weighting coefficients for the text generation loss, Binary Cross-Entropy loss, Dice loss, Semantic Segmentation loss, and Change Similarity loss, respectively. $\mathcal{L}_{txt}$ is the standard autoregressive Cross-Entropy Loss applied to the LMM’s generated sequence. $\mathcal{L}_{BCE}$ is the Binary Cross-Entropy Loss, which optimizes the model’s prediction confidence by penalizing differences between the predicted probability map and the binary ground truth at the pixel level. $\mathcal{L}_{Dice}$ is the Dice Loss, which focuses on evaluating the overlap between the predicted mask and the ground truth mask. It is particularly effective for mitigating the influence of class imbalance inherent in change detection tasks. $\mathcal{L}_{SS}$ is the 2D Cross-Entropy Loss for semantic segmentation, which computes pixel-wise classification errors across all spatial locations in the dual-temporal predictions, thereby penalizing misclassification across all semantic categories. $\mathcal{L}_{CS}$ is the Change Similarity Loss, which employs cosine embedding to constrain the semantic feature representations across the two time points, encouraging similarity in unchanged areas and divergence in changed areas.
Crucially, during training on BCD datasets, the semantic constraint losses, $\mathcal{L}_{SS}$ and $\mathcal{L}_{CS}$, are set to zero. They are only computed when training on fully annotated SCD datasets. By optimizing this comprehensive objective function, UniChange ensures effective alignment between linguistic instructions and pixel-level perception, simultaneously achieving accuracy in both binary and multi-class semantic change detection.
@@ 4.1 Experimental Setting
We employ LLaVA-7B-v1-1 [29] as the base MLMM. For the Change-Perception Vision Module, we adopt the RSBuilding-ViT-L [40] backbone, which is a SAM backbone structure pre-trained on remote sensing imagery.
We adopt 4 NVIDIA 80G H100 GPUs for training, and the training scripts leverage the deepspeed [34] engine. We utilize the AdamW [30] optimizer and set the base learning rate to 5 $\times$ 10-5. We set the batch size per device to $1$ and the gradient accumulation step to $8$. We train the model for 10 epochs, and each epoch consists of 400 training steps.
The comprehensive loss function, which was introduced in Sec. 3.3, utilizes several weighting coefficients to balance its multiple objective components. Specifically, we set the MLMM’s text generation loss weight $\lambda_{txt}$ to $1.0$. The mask loss components are weighted with $\lambda_{BCE}$ set to $2.0$, $\lambda_{Dice}$ set to $0.5$, $\lambda_{SS}$ set to $0.5$, and $\lambda_{CS}$ set to $1.0$.
4.2 Datasets
BCD Datasets. We evaluate the Binary Change Detection (BCD) capability of UniChange on three widely adopted remote sensing datasets: WHU-CD [24], LEVIR-CD+ [3], and S2Looking [37]. The WHU-CD dataset consists of a pair of large dual-temporal remote sensing images with a resolution of 0.075 m/pixel, each with a size of 32507 $\times$ 15354 pixels. For standardized evaluation, we crop these large images into 1024 $\times$ 1024 patches. We then split these patches into training, validation, and test sets following an 8:1:1 ratio. The LEVIR-CD+ dataset is an extended version of LEVIR-CD, offering 985 pairs of 1024 $\times$ 1024 images primarily focusing on building changes. We utilize the original 1024 $\times$ 1024 image size for training and testing on this dataset, which provides 637 training patches and 348 testing patches. Finally, the S2Looking dataset provides a total of 5000 image pairs captured from side-looking rural areas globally, featuring a resolution of 0.5 to 0.8 m/pixel and containing over 65920 annotated change instances. These image pairs are divided into 3500 for training, 1000 for testing and 500 for validation. These datasets collectively allow for robust assessment across different resolutions and geographic regions.
SCD datasets. We valuate the Semantic Change Detection (SCD) capabilities of UniChange on the SECOND dataset [44]. This dataset is specifically designed to support fine-grained change analysis, requiring models to simultaneously pinpoint the location of changes and classify the semantic category transitions. The dataset comprises 4662 pairs of dual-temporal remote sensing images with a 512 $\times$ 512 resolution, featuring a ground sampling distance ranging from 0.3 m to 0.5 m. It covers six detailed land-cover categories: buildings, low vegetation, trees, water, playgrounds, and bare ground. The predominance of unchanged pixels in the dataset creates a significant class imbalance, which places higher demands on the model’s generalization capability.
4.3 Evaluation Metrics
BCD Metrics. To evaluate the performance of UniChange on the BCD task, we employ four standard pixel-level metrics: Precision (P), Recall (R), F1 Score (F1), and Intersection over Union (IoU). Precision measures the proportion of correctly predicted change pixels among all pixels classified as change. Recall indicates the proportion of true positive pixels among all truly positive pixels in the ground truth. The F1 Score is the harmonic mean of Precision and Recall, providing a single metric that balances these two measures. Finally, IoU measures the overlap between predicted and ground-truth positive regions, serving as a robust measure of segmentation quality.
P $=\text{TP / (TP + FP)},$ (13)
R $=\text{TP / (TP + FN)},$
F1 $=\text{2}\times\text{P}\times\text{R}/(\text{P}+\text{R}),$
IoU $=\text{TP / (TP + FP + FN)}.$
SCD Metrics. The assessment of SCD model performance is executed using a collection of specialized metrics, all derived from the confusion matrix $Q=q_{ij}$, where $q_{ij}$ records the count of pixels classified as class $i$ with a ground truth label of $j$.
The mean Intersection over Union for SCD (mIoU) is utilized to evaluate overall segmentation quality, established as the arithmetic mean of the Intersection over Union for the regions without change ($\text{IoU}_{nc}$) and all changing regions ($\text{IoU}_{c}$):
$\text{mIoU}=(\text{IoU}_{nc}+\text{IoU}_{c})/2,$ (14)
the term $\text{IoU}_{nc}$ measures the overlap between the predicted unchanged regions and the ground-truth unchanged regions:
$\text{IoU}_{nc}=q_{00}/(\sum_{i=0}^{N}q_{i0}+\sum_{j=0}^{N}q_{0j}-q_{00}),$ (15)
@@ 4.3 Evaluation Metrics
FC-Siam-diff [9] 48.84 88.96 63.06 46.05 83.49 32.32 46.60 30.38 80.88 77.65 79.23 65.60
DASNet [5] 83.77 91.02 87.24 77.37 45.06 48.71 47.29 30.97 77.51 78.03 77.77 63.63
SNUNet [15] 91.28 87.25 89.22 80.54 45.25 50.60 47.78 31.39 78.90 78.23 78.56 64.69
BIT [4] 91.56 87.84 89.66 81.26 70.26 56.53 62.65 45.61 82.37 79.73 81.03 68.11
ChangeFormer [2] 91.76 84.85 88.17 78.85 70.73 49.25 58.07 40.91 84.29 80.81 82.51 70.23
RFL-CDNet [18] 91.33 91.46 91.39 84.15 65.72 60.82 63.17 46.17 79.95 84.04 81.94 69.41
SAM-CD [12] 96.87 85.67 90.92 83.35 72.80 58.92 65.13 48.29 79.71 81.35 81.96 69.43
Meta-CD [20] 89.00 90.35 89.67 81.27 74.08 54.00 62.47 45.42 80.17 84.13 82.10 69.93
SA-CDNet [19] 95.29 93.67 94.47 89.52 81.28 56.24 66.48 49.79 85.55 83.44 84.43 73.06
ChangeCLIP [14] 96.02 93.58 94.78 90.08 - - - - 88.46 83.90 86.12 75.63
SFCD-Net [48] 95.42 93.14 94.27 89.16 72.60 61.70 66.71 50.05 87.25 85.65 86.44 76.12
TTP [6] 96.05 92.76 94.37 89.34 73.51 62.19 67.38 50.80 85.81 84.36 85.08 74.03
@@
The Separation kappa coefficient (SeK) provides a valuable measure of semantic discrimination amidst class imbalance, particularly designed to diminish the influence of the prevalent unchanged class. SeK is computed from the confusion matrix $Q=q_{ij}$, where $\hat{q}_{ij}=q_{ij}$, but $\hat{q}_{00}=0$, and is calculated as:
The performance of UniChange on the SCD task is comprehensively evaluated using BCD metrics (IoU and F1) and SCD metrics (mIoU and SeK).
@@ 4.4 Results
BCD Results. As shown in Tab. 1, UniChange achieves the best F1 score and IoU score across all three evaluated benchmarks: WHU-CD, S2Looking, and LEVIR-CD+. On the WHU-CD dataset, our method achieves a top F1 score of 94.96 and an IoU score of 90.41. On the S2Looking dataset, our method obtains an F1 score of 69.32 and an IoU score of 53.04, exceeding the TTP result (67.38 F1 score, 50.80 IoU score). Furthermore, on the LEVIR-CD+ dataset, our approach demonstrates superior performance, setting a new SOTA F1 score of 88.19 and IoU score of 78.87, outperforming the second-best model, SFCD-Net (86.44 F1 score, 76.12 IoU score). The ability to achieve consistently superior F1 score and IoU score across these diverse benchmarks highlights the effectiveness and broad applicability of UniChange.
Table 2: Comparison with existing sota methods on SECOND.
Methods IoU mIoU F1 SeK
HRSCD.Str3 [9] 44.52 63.57 61.61 8.89
HRSCD.Str4 [9] 53.34 69.44 69.57 15.97
SCDNet [32] - 70.78 - 17.23
ChangeMask [52] 52.17 69.67 68.57 14.75
SAAN [21] 53.49 - - 18.03
BiSRNet [11] 55.99 72.38 71.79 19.64
DFINet [39] - 72.61 - 20.12
MLFA-Net [13] 56.33 72.45 72.06 20.11
MTSCD-Net [7] 55.84 71.68 71.67 20.57
SAM-SCD [31] 56.82 71.92 72.46 20.60
UniChange 57.62 72.85 73.12 23.02
SCD Results. As shown in Tab. 2, UniChange achieves the best performance across all key metrics on the SECOND dataset. UniChange achieves the highest IoU score of 57.62 and the top mIoU score of 72.85, confirming its superior capability in accurately delineating the overall regions of change and non-change. Furthermore, the model sets the SOTA F1 score at 73.12, surpassing SAM-SCD (72.46). Most notably, on the highly rigorous SeK metric, which provides a robust measure of overall semantic accuracy and inter-class differentiation, our approach obtains the highest value of 23.02, outperforming the next best performance from SAM-SCD (20.60). This comprehensive superiority across all metrics confirms that UniChange’s architecture is highly effective for handling the challenging task of multi-class change detection.
@@ 4.5 Ablation Study
In this section, we aim to validate the efficacy of the proposed components and strategies by conducting a series of ablation experiments. Unless otherwise specified, we employ LLaVA-7B-v1-1 as the MLMM and RSBuilding-ViT-L as our vision backbone. Furthermore, semantic supervision is applied to the dual-temporal remote sensing images within the WHU-CD dataset. The lora rank is set to 8, lora alpha is consistently set to twice the lora rank.
Effects of dual-temporal images Semantic Supervision. The WHU-CD dataset provides comprehensive annotations, offering not only the standard change mask but also individual semantic segmentation masks for each of the dual-temporal remote sensing images ($\text{T}_{1}$ and $\text{T}_{2}$). To validate the effects of this dual-temporal semantic supervision, we conducted an ablation study categorizing experiments into four groups. As shown in Tab. 3, The optimal performance is realized by the final group, which applied semantic supervision to both temporal images ($\text{T}_{1}$ and $\text{T}_{2}$). The intermediate configurations (supervising only $\text{T}_{1}$ or $\text{T}_{2}$) achieve progressive performance improvements over the baseline (which lacks semantic supervision), but are outperformed by the fully supervised setting. This monotonic trend confirms the synergistic effect of dual-temporal semantic supervision in enhancing feature discrimination, thereby affirming its efficacy for the change detection task within our unified framework.
Table 3: Effects of dual-temporal image semantic supervision.
@@
Effects of Different Vision Backbone. To validate the effects of different vision backbones on performance, we conduct a series of experiments as shown in Tab. 4. We test three distinct encoders (SAM, SAM2, and RSBuilding-ViT-L) under two configurations: a frozen setting and a fine-tuned (ft) setting. The SAM and RSBuilding backbones have the exact same architecture, differing only in their pre-trained weights. The experimental results reveal two clear trends. First, the fine-tuned group’s overall performance is superior to the frozen group’s across both datasets. Second, RSBuilding-ViT-L delivers the highest performance within both the fine-tuned and frozen groups.
Table 4: Effects of different vision backbone.
@@
SAM [25] 14.63 67.78 59.51 42.36
SAM2 [35] 13.84 67.48 59.13 41.97
RSBuilding-ViT-L [40] 19.30 71.11 66.76 50.10
@@
Effects of Lora Rank. To validate the effects of different LoRA ranks on model performance, we conduct experiments with ranks of 4, 8, 16, and 32, with the results detailed in Tab. 5. As shown in the table, the model’s performance varies with different ranks, achieving optimal results when the LoRA rank is set to 8.
Table 5: Effects of different lora rank.
@@
Effects of Joint Training Datasets. To validate the effects of joint training with multiple datasets, we design a progressive experimental setup, with results presented in Tab. 6. In this study, A, B, C, and D represent the WHU-CD, S2Looking, LEVIR-CD+, and SECOND datasets, respectively. We progressively increase the number of datasets in the training mixture, starting with only A, then A+B, followed by A+B+C, and finally A+B+C+D. A trend emerges from the results: as more datasets are added to the training mixture, the model’s overall performance shows an upward trend. This indicates that joint training allows the model to learn more general and robust representations, leading to improved overall performance. The best results for all datasets are achieved when the model is trained jointly on all four datasets.
Table 6: Effects of joint training datasets. A, B, C, and D represent the WHU-CD, S2Looking, LEVIR-CD+, and SECOND datasets.

### v2

@@ 3.3 Training Strategy
Our method is trained end-to-end to jointly optimise UniChange. The overall objective function, $\mathcal{L}_{total}$, which guides the entire optimisation process, is defined as a summation of all contributing loss terms:
$\mathcal{L}_{total}=\mathcal{L}_{txt}+\mathcal{L}_{mask}.$ (12)
In Eq. 12, $\mathcal{L}_{txt}$ is the standard autoregressive Cross-Entropy loss applied to the MLMM’s generated token sequence, optimising its next-token prediction capability.
The comprehensive mask loss, $\mathcal{L}_{mask}$, is designed to enforce pixel-level accuracy across the unified change detection task. This loss is defined as the summation of four distinct components:
$\mathcal{L}_{mask}=\lambda_{BCE}\mathcal{L}_{BCE}+\lambda_{Dice}\mathcal{L}_{Dice}+\lambda_{SS}\mathcal{L}_{SS}+\lambda_{SC}\mathcal{L}_{SC}.$
In Eq. 13, $\lambda_{BCE}$, $\lambda_{Dice}$, $\lambda_{SS}$, and $\lambda_{SC}$ serve as the weighting coefficients for the Binary Cross-Entropy loss, Dice loss, Semantic Segmentation loss, and Semantic Change loss, respectively. $\mathcal{L}_{BCE}$ and $\mathcal{L}_{Dice}$ are the Binary Cross-Entropy loss and the Dice loss, respectively. Both of them optimise the model’s prediction by applying supervision between the predicted change masks and the ground truth change masks. $\mathcal{L}_{SS}$ is the Cross-Entropy loss for Semantic Segmentation, which computes pixel-wise classification errors across all spatial locations between the dual-temporal predictions and their respective ground truth labels, thereby penalizing misclassification across all semantic categories. $\mathcal{L}_{SC}$ is the Semantic Change loss, designed to enhance the discriminability of dual-temporal feature representations. This loss calculates a cosine embedding distance between the semantic feature maps based on the binary ground truth masks. Specifically, it enforces similarity for unchanged regions and encourages divergence for changed regions.
Crucially, during training on BCD datasets, the semantic constraint losses, $\mathcal{L}_{SS}$ and $\mathcal{L}_{SC}$, are set to zero. They are only computed when training on SCD datasets. By optimising this comprehensive objective function, UniChange ensures both robust and effective alignment between textual instructions and pixel-level perception, simultaneously achieving high accuracy in both binary and multi-class semantic change detection tasks.
@@ 4.1 Experimental Setting
We employ LLaVA-7B-v1-1 [liu2023visual] as the base MLMM. For Vision Backbone, we adopt the RSBuilding-ViT-L [wang2024rsbuilding] backbone, which is a SAM backbone structure pre-trained on remote sensing imagery.
UniChange is trained for 10 epochs (with 400 steps per epoch) using 4 NVIDIA 80G H100 GPUs. We employ the AdamW [loshchilov2017decoupled] optimizer with a base learning rate of $5\times 10^{-5}$. The training is performed using a per-device batch size of 1 and a gradient accumulation step of 8, leveraging the deepspeed [rasley2020deepspeed] engine for efficiency.
The comprehensive loss function, which is introduced in Sec. 3.3, utilises several weighting coefficients to balance its multiple objective components. Specifically, the mask loss components are weighted with $\lambda_{BCE}$ set to $2.0$, $\lambda_{Dice}$ set to $0.5$, $\lambda_{SS}$ set to $0.5$, and $\lambda_{SC}$ set to 1.0.
@@
FC-Siam-diff [daudt2018fully] 48.84 88.96 63.06 46.05 83.49 32.32 46.60 30.38 80.88 77.65 79.23 65.60
DASNet [chen2020dasnet] 83.77 91.02 87.24 77.37 45.06 48.71 47.29 30.97 77.51 78.03 77.77 63.63
SNUNet [fang2021snunet] 91.28 87.25 89.22 80.54 45.25 50.60 47.78 31.39 78.90 78.23 78.56 64.69
BIT [chen2022remote] 91.56 87.84 89.66 81.26 70.26 56.53 62.65 45.61 82.37 79.73 81.03 68.11
ChangeFormer [bandara2022transformer] 91.76 84.85 88.17 78.85 70.73 49.25 58.07 40.91 84.29 80.81 82.51 70.23
RFL-CDNet [gan2024rfl] 91.33 91.46 91.39 84.15 65.72 60.82 63.17 46.17 79.95 84.04 81.94 69.41
SAM-CD [ding2024adapting] 96.87 85.67 90.92 83.35 72.80 58.92 65.13 48.29 79.71 81.35 81.96 69.43
Meta-CD [gao2025combining] 89.00 90.35 89.67 81.27 74.08 54.00 62.47 45.42 80.17 84.13 82.10 69.93
SA-CDNet [gan2025detect] 95.29 93.67 94.47 89.52 81.28 56.24 66.48 49.79 85.55 83.44 84.43 73.06
ChangeCLIP [dong2024changeclip] 96.02 93.58 94.78 90.08 - - - - 88.46 83.90 86.12 75.63
Changer [fang2023changer] - - - - 73.01 62.04 67.08 50.47 - - - -
LSKNet [li2025lsknet] - - - - 71.90 63.64 67.52 50.96 - - - -
SFCD-Net [zhang2024integrating] 95.42 93.14 94.27 89.16 72.60 61.70 66.71 50.05 87.25 85.65 86.44 76.12
TTP [chen2024time] 96.05 92.76 94.37 89.34 73.51 62.19 67.38 50.80 85.81 84.36 85.08 74.03
@@
4.2 Datasets
BCD Datasets. We evaluate the binary change detection (BCD) capability of UniChange on three widely adopted remote sensing datasets: WHU-CD [ji2019fully], S2Looking [shen2021s2looking], and LEVIR-CD+ [chen2020spatial]. The WHU-CD dataset consists of a pair of large dual-temporal remote sensing images, each with a size of 32507 $\times$ 15354. For standardised evaluation, we crop these large images into 1024 $\times$ 1024 patches. We then split these patches into training, validation, and testing sets following an 8:1:1 ratio. The S2Looking dataset provides a total of 5000 image pairs captured from side-looking rural areas globally. These image pairs are divided into 3500 for training, 1000 for testing and 500 for validation. Finally, LEVIR-CD+ is an extended version of LEVIR-CD. It provides 985 pairs of 1024 $\times$ 1024 images, primarily focusing on building changes. We utilise the original 1024 $\times$ 1024 image size for training and testing, which provides 637 training patches and 348 testing patches. These datasets collectively allow for robust assessment across different resolutions and geographic regions.
SCD Datasets. We evaluate the semantic change detection (SCD) capability of UniChange on the SECOND dataset [yang2022asymmetric]. The dataset is specifically designed for fine-grained change analysis, requiring models to simultaneously localise changes and classify semantic transitions. The dataset comprises 4662 pairs of dual-temporal remote sensing images with a 512 $\times$ 512 resolution. It covers six detailed land-cover categories: building, low vegetation, tree, water, playground, and bare ground. The predominance of unchanged pixels in the dataset creates a significant class imbalance, which places higher demands on the model’s generalisation capability.
4.3 Evaluation Metrics
BCD Metrics. To evaluate BCD performance, we employ four standard pixel-level metrics: Precision (P), Recall (R), F1 score (F1), and Intersection over Union (IoU). Precision measures the ratio of correctly predicted change pixels to all predicted change pixels. Recall measures the ratio of true positive pixels to all ground-truth positive pixels. The F1 score is the harmonic mean of P and R. IoU quantifies the overlap between predicted and ground-truth change regions.
SCD Metrics. To evaluate SCD performance, we employ three metrics: mIoU, $\text{F}_{scd}$, and SeK. mIoU evaluates the overall segmentation quality. $\text{F}_{scd}$ quantifies semantic transition accuracy using the harmonic mean of precision and recall. SeK measures semantic discrimination, mitigating the impact of the unchanged class. UniChange’s SCD performance is comprehensively evaluated using BCD metrics (IoU, F1 ($\text{F}_{bcd}$)) and SCD metrics (mIoU, $\text{F}_{scd}$, SeK). Further details on the BCD and SCD metrics are provided in Appendix A of the supplementary materials.
@@ 4.4 Results
BCD Results. As shown in Tab. 1, UniChange achieves the best F1 score and IoU score across all three evaluated benchmarks: WHU-CD, S2Looking, and LEVIR-CD+. On the WHU-CD dataset, our method achieves top F1 and IoU scores of 94.96 and 90.41. On the S2Looking dataset, our method obtains an F1 score of 69.32 and an IoU score of 53.04, exceeding the LSKNet result (67.52 F1 score, 50.96 IoU score). Furthermore, on the LEVIR-CD+ dataset, our approach demonstrates superior performance, setting a new SOTA F1 score of 88.19 and IoU score of 78.87, outperforming the second-best model, SFCD-Net (86.44 F1 score, 76.12 IoU score). This consistent superiority in F1 and IoU scores across diverse benchmarks highlights UniChange’s effectiveness and broad applicability.
Table 2: Comparison with existing SOTA methods on SECOND. * denotes re-implementation under identical settings.
Methods IoU mIoU $\text{F}_{scd}$ $\text{F}_{bcd}$ SeK
HRSCD.Str4 [daudt2019multitask] 53.34 69.44 - 69.57 15.97
ChangeMask [zheng2022changemask] 54.23 - - 70.32 17.89
SAAN [guo2024saan] 53.49 - - - 18.03
DESNet [wang2023difference] - 70.73 58.75 - 17.97
STSP-Net [he2023spatial] - 72.03 60.77 72.05 20.91
DFINet [wang2024dual] - 72.61 - - 20.12
MLFA-Net [ding2024mlfa] 56.33 72.45 - 72.06 20.11
TextSCD [huang2025textscd] - 72.38 61.90 - 21.66
GLAI-Net [ding2025glai] 56.21 - - 71.97 20.63
Bi-SRNet* [ding2022bi] 56.89 72.25 62.12 72.52 21.69
SCDNet* [peng2021scdnet] 55.51 71.63 62.40 71.39 21.43
MTSCD-Net* [cui2023mtscd] 54.10 68.99 58.31 70.21 18.83
SCD-SAM* [mei2024scd] 56.82 71.92 60.71 72.46 20.60
HGINet* [long2024semantic] 56.13 71.73 62.88 71.90 21.83
MambaSCD* [chen2024changemamba] 57.24 72.73 54.17 72.81 21.31
UniChange 57.62 72.85 63.50 73.12 23.02
SCD Results. As shown in Tab. 2, UniChange achieves the highest IoU score, mIoU score and $\text{F}_{bcd}$ score, confirming its superior capability in accurately delineating the overall regions of change and non-change. Furthermore, the model’s SOTA performance on $\text{F}_{scd}$ and SeK metrics verifies that UniChange’s semantic change perception capability is equally outstanding. This dual excellence across both binary and semantic accuracy capabilities exemplifies the unique, robust unification achieved by our MLLM-based framework, setting it apart from prior specialised models. For instance, MambaSCD, while achieving a strong binary detection result (mIoU 72.73, $\text{F}_{bcd}$ 72.81), exhibits a substantial drop in semantic accuracy ($\text{F}_{scd}$ 54.17). Similarly, HGINet, despite attaining a competitive semantic score ($\text{F}_{scd}$ 62.88, the second highest), shows weaker binary detection performance (mIoU 71.73, $\text{F}_{bcd}$ 71.90). This comprehensive superiority across all metrics confirms that UniChange’s architecture is highly effective for handling the challenging task of multi-class change detection.
@@ 4.5 Ablation Study
In this section, we aim to validate the efficacy of the proposed components and strategies by conducting a series of ablation experiments. Unless otherwise specified, we employ LLaVA-7B-v1-1 as the base MLMM and RSBuilding-ViT-L as our vision backbone. Furthermore, semantic supervision is applied to the dual-temporal remote sensing images within the WHU-CD dataset. The lora rank is set to 8, and lora alpha is consistently set to twice the lora rank.
Effects of Dual-Temporal Images Semantic Supervision. The WHU-CD dataset provides comprehensive annotations, including a change mask and individual semantic segmentation masks for both dual-temporal images ($\text{T}_{1}$ and $\text{T}_{2}$). To validate the effects of this dual-temporal semantic supervision, we conduct an ablation study categorizing experiments into four groups. As shown in Tab. 3, the optimal performance is realized by the final group, which applies semantic supervision to both temporal images ($\text{T}_{1}$ and $\text{T}_{2}$). The intermediate configurations (supervising only $\text{T}_{1}$ or $\text{T}_{2}$) achieve progressive performance improvements over the baseline (lacking semantic supervision), but are outperformed by the fully supervised setting. This monotonic trend confirms the synergistic effect of dual-temporal semantic supervision in enhancing feature discrimination, thereby affirming its efficacy within our unified framework.
Table 3: Effects of dual-temporal images semantic supervision.
@@
Effects of Different Vision Backbones. To validate the effects of different vision backbones, we conduct experiments using three backbones (SAM, SAM2, and RSBuilding-ViT-L) under frozen and fine-tuned (ft) configurations. Note that SAM and RSBuilding backbones share the same architecture, differing only in pre-trained weights. Results in Tab. 4 show two trends: First, the fine-tuned group consistently outperforms the frozen group. Second, RSBuilding-ViT-L delivers best results in both settings.
Table 4: Effects of different vision backbones.
@@
SAM [kirillov2023segment] 14.63 67.78 59.51 42.36
SAM2 [ravi2024sam] 13.84 67.48 59.13 41.97
RSBuilding-ViT-L [wang2024rsbuilding] 19.30 71.11 66.76 50.10
@@
Effects of LoRA Ranks. To validate the effects of different LoRA ranks on model performance, we conduct experiments with LoRA ranks of 4, 8, 16, and 32, with the results detailed in Tab. 5. As shown in the table, the model’s performance varies with different LoRA ranks, achieving optimal results when the LoRA rank is set to 8.
Table 5: Effects of different LoRA ranks.
@@
Effects of Joint Training Datasets. We design a progressive experimental setup to validate the effects of joint training with multiple datasets, with results presented in Tab. 6. In this study, A, B, C, and D represent the WHU-CD, S2Looking, LEVIR-CD+, and SECOND datasets, respectively. We progressively increase the number of datasets in the training mixture, starting with A, then A+B, A+B+C, and finally A+B+C+D. Results show a clear upward trend in overall performance as more datasets are added. This indicates that joint training allows the model to learn more general and robust representations. The best results for all datasets are achieved when the model is trained jointly on all four datasets.
Table 6: Effects of joint training datasets.